solve for a.
n= 353/172-a
please show me steps to do this problem.
Thanks

Answer by stanbon(75887)

solve for a.
n= 353/172-a
-----
I'm going to assume the denominator is (172-a).
If that is correct you get:
(172-a) = 353/n
----
Then a = 172 - [352/n]
---
Or
a = [172n-352]/n
===================
Cheers,
Stan H.
===================